OpenStudy (makayla):

42 is subtracted from 42% of a number. The result is 42. What is the number?

OpenStudy (anonymous):

sklee can u help me again

OpenStudy (anonymous):

(42/100)x - 42 = 42 (42/100)x = 84 x = 84(100/42) = 200